OpenStudy (anonymous):

Input the equation of the given line in slope-intercept form. The line with m = 2, intersecting 2x - 3y = 6 at x = 3.

OpenStudy (anonymous):

slope intercept form: 3y=2x-6 y=2/3 x -2

OpenStudy (anonymous):

the line with m=2 @ x=3 y-y1=2(x-3) you need to find y1. because 2 line intersecting @ x=3 , y coordinate is the same as well. let's find it: y=2/3 *3 -2=0 so point of intersection is: (3,0) the equation of the desired line will be: y-0=2(x-3) y=2x-6